OpenStudy (anonymous):

the mass of a regular solid is 160 g. When measured with a metric ruler each side of the cube measures 2cm. What is the density of that object?

OpenStudy (anonymous):

Density= Mass/ Volume Here mass = 160 g Volume = 2^3= 8 So, 160g/8cm^3=?